solve the equation in exact differentials

(y2-1)dx+(2xy+3y)dy=0


1
Expert's answer
2022-01-07T12:31:06-0500

Let us solve the equation in exact differentials "(y^2-1)dx+(2xy+3y)dy=0,"

Since "\frac{\partial (y^2-1)}{\partial y}=2y=\frac{\partial (2xy+3y)}{\partial x}", the differential equation is indeed exact.

It follows that there exists the function "u=u(x,y)" such that

"\frac{\partial u}{\partial x}=y^2-1,\ \frac{\partial u}{\partial y}=2xy+3y." Therefore, "u(x,y)=xy^2-x+c(y)," and hence

"\frac{\partial u}{\partial y}=2xy+c'(y)=2xy+3y." It follows that "c'(y)=3y," and thus "c(y)=\frac{3}2y^2+C."

We conclude that the general solution is

"xy^2-x+\frac{3}2y^2=C."